OpenStudy (anonymous):

what is the midpoint for(9,9)(4,0)

OpenStudy (aravindg):

use midpoint formula \[\huge Midpoint = (\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2})\]

OpenStudy (anonymous):

xp=(9+4)/2 yp=(9)/2 mp(13/2,4.5)

OpenStudy (aravindg):

srry its 4.5

OpenStudy (anonymous):

thanks

OpenStudy (aravindg):

midpoint=(13/2,4.5)